OpenStudy (anonymous):

Find a positive coterminal angle to 36 degrees. A. 120 degrees B. 270 degrees C. 366 degrees D. 396 degrees

OpenStudy (anonymous):

First add 360 to 36 to get your positive coterminal angle. 36+360=396 degrees

OpenStudy (anonymous):

Yikes, sorry, forget that last one..let me redo these. Was looking at the wrong question.

OpenStudy (anonymous):

To get the positive coterminal angle, you have to add 36 + 360 =? To get the the negative coterminal, you have to subtract 36 - 360=?

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

Obviously, you're asking for the positive coterminal, so you use the first option.